OpenStudy (anonymous):

The greatest common divisor of positive integers m and n is 8. The least common multiple of m and n is 112. What is the least possible value of m + n?

OpenStudy (anonymous):

What I thought was that the l.c.m. multiplied by the g.c.d. of two numbers is equal to m*n, when I multiplied that out I got 869, did prime factorization and got 128*7.... where am I wrong?

OpenStudy (anonymous):

hello?

OpenStudy (rizags):

hello?

OpenStudy (rational):

m = 8m' n = 8n' mn = 8*112 m'n' = 14 so (m', n') = (1, 14), (2, 7), (7, 7) the least value of m' + n' is 2+7 = 9 so the least value of m+n is 8*9 = 72

OpenStudy (anonymous):

but, what is m' and n' ? what is the meaning of ( ' )?

OpenStudy (rizags):

lemme do a bit of casework on this one

OpenStudy (rizags):

alrite, so to begin, we know that \[m, n \le 112\] because their least common multiple is 112, and multiples of positive integers are always greater than or equal to themselves

OpenStudy (rizags):

yea that guy rational up there, is right

OpenStudy (rizags):

his solution is very elegant, but heres a simpler one

OpenStudy (rizags):

to begin, list all factors of 112 in ascending order for me

OpenStudy (rizags):

heloooooo? could you do that ^^